# Higashi-Hiroshima, Hiroshima 739-8526, Japan

### v1

@@ Abstract
We propose a bootstrap approximation method for the Hermitian one-matrix model that does not rely on positivity constraints. The theoretical foundation of this method is that the one-matrix model admits an eigenvalue distribution $\rho(\lambda)$, and that the moments $w_{n}$ generated from it satisfy the loop equations. Our framework is designed to numerically determine a self-consistent pair of $\rho(\lambda)$ and $w_{n}$ that simultaneously satisfies these two requirements. In the concrete implementation, we employ a least-squares method, for which no sign problem arises in principle, and therefore the method can be formally applied also to Minkowski-type models. Actual numerical calculations show that this bootstrap approximation reproduces, with very high accuracy, the exact solutions for Euclidean-type models and the perturbative results for Minkowski-type models.
@@ 1 Introduction
In this work, we propose a numerical method for large-$N$ Minkowski-type matrix models. Unlike the existing matrix bootstrap approach, the present method does not use positivity at all. Instead, we employ the eigenvalue distribution $\rho(\lambda)$, which is characteristic of large-$N$ matrix models, and impose self-consistency constraints between this distribution and the moments $w_{n}=\langle\operatorname{tr}\phi^{n}\rangle$. In this formulation, the constraints amount to requiring $|w_{n}-w_{n}^{(P)}|\simeq 0$. Since this quantity is evaluated in terms of its absolute value, numerical computations can be carried out without being affected by sign problems or by the breakdown of positivity.
The structure of this paper is as follows. In Section 2, where the method is introduced, we begin with a brief review of the Hermitian one-matrix model that will serve as a toy model in this work, and then discuss positivity, which forms the theoretical foundation of usual bootstrap approaches. In contrast to this, we introduce the eigenvalue distribution as one of the central concepts of the present study and explain how it can be approximated by polynomials. In Section 3, we discuss how this method can be extended to Minkowski-type theories. Section 4 is devoted to concrete numerical calculations, where we demonstrate that known results for the one-matrix model are successfully reproduced in both the Euclidean and Minkowski cases. Finally, in Section 5, we summarize the method and the numerical results, and present several perspectives on how the approach can be extended to more realistic models, such as the IKKT matrix model and the (twisted) Eguchi–Kawai model. Some technical but important calculations and facts related to the one-matrix model are collected in the Appendices.
@@ 2.1 Matrix bootstrap with positivity constraint
In view of the limitations of Monte Carlo methods—namely, their restriction to finite $N$ and the sign problem—the central requirement we impose on a matrix-bootstrap approach is the following: it should be capable of analyzing large-$N$ Minkowski-type matrix models. The main result of the present work is to demonstrate that such a bootstrap is indeed possible by exploiting the eigenvalue distribution of the one-matrix model.
@@ 2.2 Eigenvalue distribution
$\rho_{E}(\lambda)\equiv\left\langle\frac{1}{N}\sum_{k=1}^{N}\delta(\lambda-\lambda_{k})\right\rangle_{E}=\left\langle\operatorname{tr}\!\left(\frac{1}{\lambda-\phi}\right)\right\rangle_{E}.$ (11)
@@ 2.3.2 Theoretical basis of the formulation
In this paper, we do not pursue the issue of non-uniqueness any further. Instead, because our procedure is explicitly based on the existence of the eigenvalue density $\rho_{E}(\lambda)$, we regard the resulting approximate values as strong candidates for physical solutions. At the same time, it should be kept in mind that, independently of the issue of non-uniqueness, the present method is intrinsically approximate. In order to justify the resulting approximate values as candidates for physical solutions, it is first necessary to verify that this approximation indeed provides a (algebraic) solution to the loop equations, as will be discussed in Section 2.4.
@@ 3.1 Large-$N$ master field: a “density-matrix” interpretation
Finally, all the arguments presented above rely on the assumptions made in (42), (43), and (44), and it should be emphasized that our aim is not to justify these assumptions themselves. Rather, our aim is to render the bootstrap approximation applicable to Minkowski-type matrix models, and the assumptions above should be regarded merely as hypotheses that make numerical analysis feasible. The success or failure of this approach is then judged by whether the objective function $F$ and the test function $G$ can be made sufficiently small, that is, whether the existence of nontrivial (algebraic) solutions to the loop equations is suggested under these assumptions.
@@ 3.3 Regularization
Turning to the treatment of the Minkowski-type one-matrix model in the present work, no explicit regularization based on the $i\epsilon$ prescription has been carried out so far. This is because our stance in this study is to find consistent solutions of the loop equations (5) and (51), and to demonstrate that such solutions are physical, i.e., consistent with the existence of a (formal) eigenvalue distribution. Within this framework, it is not necessary to explicitly invoke the matrix integral itself. In addition, as seen in Section 4, the present bootstrap approximation for the Minkowski-type one-matrix model reproduces perturbative results well, and moreover appears to provide a good fit to the formal eigenvalue distribution $\rho_{M}(\lambda)$.
@@ 4.2.1 Endpoint-free approximation
Since the $\mathbb{Z}_{2}$ symmetry implies $w_{2n-1}=0$, it is necessary in numerical computations to skip the odd-order moments in advance in this manner. As explained in Section 2.4, for this text function $G_{E}$ we define $L\equiv\Lambda^{\prime}-\Lambda$ using the largest $\Lambda^{\prime}$ that satisfies $G_{E}\leq 1$. A larger value of $L$ can be interpreted as an indication that the approximation is performing well.
@@ 4.2.2 Endpoint-fixed approximation
With this choice, the condition $\rho_{E}^{(P)}(\pm a)=0$ is enforced by construction. It should be noted that, compared with the endpoint-unfixed case, this ansatz reduces the number of independent coefficients $c_{2m}$ by one. Consequently, even when considering a polynomial of total degree ten, the number of unknown variables is reduced from eight to seven. In accordance with this reduction, we also decrease the number of imposed equations by setting $\Lambda=6$.
@@ 4.3 Minkowski-type one-matrix model
As discussed in Section 4.1, in the Minkowski-type one-matrix model an additional angular (phase) parameter $\theta$ appears, and the coefficients $c_{2m}$ as well as the moments $w_{2n}$ become complex-valued. As a result, the number of variables increases, and even for the same value of $M$, a larger number of loop equations is required. In the following, we adopt the endpoint-fixed ansatz (56) from the outset and, as in the Euclidean case, assume a polynomial of degree ten. In this setup, the unknown variables are $c_{0},\ldots,c_{8},\theta,a,$ and $w_{2}$, which amount to $14$ real degrees of freedom. We therefore set $\Lambda=12$ and incorporate the conditions $w_{0}=w_{0}^{(P)},w_{2}=w_{2}^{(P)},\ldots,w_{12}=w_{12}^{(P)}$ into the least-squares procedure. In this way, the number of equations matches the number of unknowns. Note that both $w_{n}$ and $w_{n}^{(P)}$ are complex, and each condition $w_{n}=w_{n}^{(P)}$ thus provides two real equations. Unlike the Euclidean model, we set $r_{n}=(0.8)^{n}$ in the present analysis in order to improve numerical stability.
Here, as a “formal eigenvalue distribution” of the Minkowski-type one-matrix model, we define $\rho_{M}(z)$ following the discussion in Appendix B. Our stance on this point is as follows. First, when the value of the test function $G$ is sufficiently small—especially when it is comparable to or even smaller than in the Euclidean case—the present approximation is likely to provide an algebraic solution of the loop equations. Furthermore, since the formal eigenvalue distribution $\rho_{M}(z)$ reproduces the perturbative expansion, if our approximation succeeds in reproducing this $\rho_{M}(z)$, it strongly suggests that the approximation also captures a “physical” solution of the loop equations.
@@
Table 4: Bootstrap approximations for the Minkowski-type one-matrix model performed with $M=8$, $\Lambda=12$, $r_{n}=(0.8)^{n}$, and $g=\pm 1,-0.3,-0.1$. The quantities $a^{\text{formal}}$ and $\theta^{\text{formal}}$ are computed according to the formal solution given in Appendix B.
@@
With these considerations in mind, Figure 3 presents a comparison between $\rho_{M}^{(P)}(z)$ obtained from the bootstrap approximation and the formal solution $\rho_{M}(z)$ with g=1g=-1. Both distributions are defined on line segments in the complex plane connecting $\pm e^{i\theta}a$. Here it should be emphasized that the parameters $\theta$ and $a$ are in general different for $\rho_{M}(z)$ and $\rho_{M}^{(P)}(z)$. To avoid confusion, we denote by $\theta$ and $a$ the parameters associated with $\rho_{M}(z)$, whose values are determined by (74), while $\theta^{(P)}$ and $a^{(P)}$ are those determined by the bootstrap approximation. With this notation, the supports can be written explicitly as
@@
Accordingly, $\rho_{M}(z)$ and $\rho_{M}^{(P)}(z)$ are defined on $\Gamma$ and $\Gamma^{(P)}$ as $\rho_{M}(e^{i\theta}\lambda)$ and $\rho_{M}^{(P)}(e^{i\theta^{(P)}}\lambda)$, respectively. When plotting the comparison between the two distributions, we evaluate $\rho_{M}(z)$ and $\rho_{M}^{(P)}(z)$ on their respective supports $\Gamma$ and $\Gamma^{(P)}$, and then display them on the same $\lambda$$\rho$ plane. As a consequence, the difference between the angles $\theta$ and $\theta^{(P)}$ cannot be read off directly from Figure 3 (however, as can be seen from Table 4, the numerical difference between these quantities is in fact extremely small).. The numerical values of the relevant parameters are summarized in Table 4.
@@
Finally, we comment on the phase transition of the Minkowski model. In the Euclidean model, there exists a critical value $g_{c}$ of $g$, beyond which a phase transition occurs and the theory breaks down. In the bootstrap approximation, this fact is presumably reflected in the failure of the optimization at $g=0.1$, even though the number of unknown variables is balanced with the number of equations. In contrast, in the Minkowski model, as far as we have investigated, no such behavior is observed, except for the instability around $g=0$. From these numerical results, we conjecture that in the Minkowski-type model no phase transition occurs for any $g\in\mathbb{R}$. However, in the Euclidean one-matrix model, the existence of a critical value $g_{c}$ played an essentially important role in obtaining the equivalence with Liouville theory via the continuum limit of the model. Whether a similar continuum limit exists in the Minkowski one-matrix model, and whether the conjecture that no critical value $g_{c}$ exists is truly correct, require further investigation.
@@ 5 Summary and Discussion
In numerical studies of matrix models, direct evaluations of matrix integrals using Monte Carlo methods have been extensively carried out, greatly advancing our understanding of theories primarily at finite $N$. On the other hand, within Monte Carlo approaches, taking the limit $N\to\infty$ is not possible in principle. From a theoretical point of view, however, the large-$N$ limit is expected to drastically simplify the structure of the system, while analytical calculations are in fact much more difficult at finite $N$, so this represents a rather curious inversion. Only recently has this situation begun to change, thanks to the development of matrix bootstrap methods based on semidefinite programming (SDP). Turning to our approach, we make more active use of the fact that large-$N$ matrix models are expected to possess a simple underlying structure, and translate this structure more directly into a numerical framework. One of the benefits of this approach is that the simplification due to the large-$N$ limit also extends to Minkowski-type matrix models, which makes it possible to apply the present method to the Minkowski-type one-matrix model.
Concretely, in this work we have proposed a bootstrap approximation method based on the eigenvalue distribution $\rho(\lambda)$ and applied it to both Euclidean- and Minkowski-type one-matrix models. The method is designed to determine the (polynomially approximated) eigenvalue distribution $\rho^{(P)}(\lambda)$ and the values of the moments $w_{1},w_{2}$ self-consistently so as to satisfy the following three conditions: (i) there exists a probability distribution $\rho(\lambda)$ that generates the moments $w_{n}$; (ii) $\rho(\lambda)$ is a continuous function with finite support and can be well approximated by a finite-order polynomial; and (iii) the moments $w_{n}$ obey the loop equations. In practice, this is achieved by implementing a least-squares method that minimizes the objective function $F=\sum_{n}^{\Lambda}\lvert w_{n}-w_{n}^{(P)}\rvert^{2}$ . In addition, in order to apply this method to Minkowski-type one-matrix models, we have formally extended the notion of the eigenvalue distribution by relying on the large-$N$ factorization and the large-$N$ master-field conjecture. Under this assumption, the method can be applied to both Euclidean- and Minkowski-type models on the same footing. In fact, numerical calculations demonstrate that the resulting solutions are in excellent agreement with the exact and formal ones.

### v2

@@ Abstract
We propose a bootstrap approximation method for the Hermitian one-matrix model that does not rely on positivity constraints. The theoretical foundation of this method is that the one-matrix model admits an eigenvalue distribution $\rho(\lambda)$, and that the moments $w_{n}$ generated from it satisfy the loop equations. Our framework is designed to numerically determine a self-consistent pair of $\rho(\lambda)$ and $w_{n}$ that simultaneously satisfies these two requirements. In the concrete implementation the least-squares method is employed, and since the sign problem is absent in this formulation, the method can be formally applied to the Minkowski one-matrix model as well, provided that the one-cut structure of the resolvent is assumed. Actual numerical calculations show that this bootstrap approximation reproduces, with very high accuracy, the exact solutions for Euclidean-type models and the perturbative results for Minkowski-type models.
@@ 1 Introduction
In this work, we propose a numerical method for large-$N$ Minkowski-type matrix models. Unlike the existing matrix bootstrap approach, the present method does not use positivity at all. Instead, we employ the eigenvalue distribution $\rho(\lambda)$, which is characteristic of large-$N$ matrix models, and impose self-consistency constraints between this distribution and the moments $w_{n}=\langle\operatorname{tr}\phi^{n}\rangle$. In this formulation, the constraints amount to requiring $|w_{n}-w_{n}^{(P)}|\simeq 0$. Since this quantity is evaluated in absolute value, the resulting numerical procedure is in principle free from the sign problem and from the breakdown of positivity.
However, in Minkowski-type matrix models the weight $e^{iS}$ is complex, and therefore it is not yet fully understood what form the eigenvalue distribution $\rho(\lambda)$ should take, or even whether such an object exists. For this reason, in the present work we consider the simplest matrix model, namely the one-matrix model, and moreover adopt the nontrivial assumption that the resolvent $R(z)$ has only a single cut. We then show that, at least as long as this assumption is valid, the numerical results reproduce the perturbative results with high accuracy.
The structure of this paper is as follows. In Section 2, where the method is introduced, we begin with a brief review of the Hermitian one-matrix model that will serve as a toy model in this work, and then discuss positivity, which forms the theoretical foundation of usual bootstrap approaches. In contrast to this, we introduce the eigenvalue distribution as one of the central concepts of the present study and explain how it can be approximated by polynomials. In Section 3, we discuss how this method can be extended to Minkowski-type theories. Section 4 is devoted to concrete numerical calculations, where we demonstrate that known results for the one-matrix model are successfully reproduced in both the Euclidean and Minkowski cases. Finally, in Section 5, we summarize the method and the numerical results, and present several perspectives on how the approach can be extended to more realistic models, such as the IKKT matrix model and the (twisted) Eguchi–Kawai model. Some technical but important calculations and facts related to the one-matrix model are collected in the Appendices. This includes analytical (and formal) results and discussions related to the Minkowski-type one-matrix model.
@@ 2.1 Matrix bootstrap with positivity constraint
In view of the limitations of Monte Carlo methods—namely, their restriction to finite $N$ and the sign problem—the central requirement we impose on a matrix-bootstrap approach is the following: it should be capable of analyzing large-$N$ Minkowski-type matrix models.
@@ 2.2 Eigenvalue distribution
$\rho_{E}(\lambda)\equiv\left\langle\frac{1}{N}\sum_{k=1}^{N}\delta(\lambda-\lambda_{k})\right\rangle_{E}.$ (11)
@@ 2.3.2 Theoretical basis of the formulation
In this paper, we do not pursue the issue of non-uniqueness any further. Instead, because our procedure is explicitly based on the existence of the eigenvalue density $\rho_{E}(\lambda)$, we regard the resulting approximate values as strong candidates for physical solutions, at least in the Euclidean-type one-matrix model. At the same time, it should be kept in mind that, independently of the issue of non-uniqueness, the present method is intrinsically approximate. In order to justify the resulting approximate values as candidates for physical solutions, it is first necessary to verify that this approximation indeed provides a (algebraic) solution to the loop equations, as will be discussed in Section 2.4.
@@ 3.1 Large-$N$ master field: a “density-matrix” interpretation
Finally, it should be noted that the discussion so far involves several important and nontrivial assumptions. Namely, (1) that the master field $\hat{\phi}$ exists, (2) that, as in the Euclidean case, the eigenvectors $|\lambda\rangle$ of $\hat{\phi}$ form a complete basis and at the same time are also eigenvectors of $\hat{\rho}_{M}$, and (3) that the eigenvalues of $\hat{\phi}$ are given by $e^{i\theta}\lambda$. Among these, the third condition corresponds to the assumption that the resolvent $R(z)$ of the Minkowski-type one-matrix model has a single cut in the complex plane. Under this assumption the perturbative expansion of the moment $w_{2}$ can be reproduced exactly, and therefore the assumption is expected to be valid at least in the region where the absolute value of the coupling constant $g$ is small (see the Appendix B for the detailed calculation).
However, in the region where $g$ becomes large, perturbation theory can no longer be trusted in a naive way, and the assumption of the one-cut structure loses its justification. This implies that, once the above three assumptions are imposed, the numerical computation necessarily loses a certain degree of generality. In particular, for relatively large values of $g$, where the perturbative expansion around $g=0$ is no longer expected to provide a reliable guide, even if the optimization converges successfully, it does not necessarily follow that the obtained solution corresponds to a genuinely physical one. This issue must ultimately be examined by a method that does not rely on the one-cut assumption, and a detailed investigation will be presented in a forthcoming paper.
@@ 3.3 Regularization
Turning to the treatment of the Minkowski-type one-matrix model in the present work, no explicit regularization based on the $i\epsilon$ prescription has been carried out so far. This is because our stance in this study is to find consistent solutions of the loop equations (5) and (51), and to demonstrate that such solutions are consistent with the existence of a (formal) eigenvalue distribution. Within this framework, it is not necessary to explicitly invoke the matrix integral itself. In addition, as seen in Section 4, the present bootstrap approximation for the Minkowski-type one-matrix model reproduces perturbative results well, and moreover appears to provide a good fit to the formal eigenvalue distribution $\rho_{M}(\lambda)$.
@@ 4.2.1 Endpoint-free approximation
Since the $\mathbb{Z}_{2}$ symmetry implies $w_{2n-1}=0$, it is necessary in numerical computations to skip the odd-order moments in advance in this manner. As explained in Section 2.4, for this test function $G_{E}$ we define $L\equiv\Lambda^{\prime}-\Lambda$ using the largest $\Lambda^{\prime}$ that satisfies $G_{E}\leq 1$. A larger value of $L$ can be interpreted as an indication that the approximation is performing well.
@@ 4.2.2 Endpoint-fixed approximation
With this choice, the condition $\rho_{E}^{(P)}(\pm a)=0$ is enforced by construction. It should be noted that, compared with the endpoint-unfixed case, this ansatz reduces the number of independent coefficients $c_{2m}$ by one. Consequently, even when considering a polynomial of total degree ten, the number of unknown variables is reduced from eight to seven. In accordance with this reduction, we also decrease the number of imposed equations by setting $\Lambda=12$.
@@ 4.3 Minkowski-type one-matrix model
As discussed in Section 4.1, in the Minkowski-type one-matrix model an additional angular (phase) parameter $\theta$ appears, and the coefficients $c_{2m}$ as well as the moments $w_{2n}$ become complex-valued. In the following, we adopt the endpoint-fixed ansatz (56) from the outset and, as in the Euclidean case, assume a polynomial of degree ten. In this setup, the unknown variables are $c_{0},\ldots,c_{8},\theta,a,$ and $w_{2}$, which amount to $14$ real degrees of freedom. We therefore set $\Lambda=12$ and incorporate the conditions $w_{0}=w_{0}^{(P)},w_{2}=w_{2}^{(P)},\ldots,w_{12}=w_{12}^{(P)}$ into the least-squares procedure. In this way, the number of equations matches the number of unknowns. Note that both $w_{n}$ and $w_{n}^{(P)}$ are complex, and each condition $w_{n}=w_{n}^{(P)}$ thus provides two real equations. Unlike the Euclidean model, we set $r_{n}=(0.8)^{n}$ in the present analysis in order to improve numerical stability.
Here, as a “formal eigenvalue distribution” of the Minkowski-type one-matrix model, we define $\rho_{M}(z)$ following the discussion in Appendix B. Our stance on this point is as follows. First, when the value of the test function $G$ is sufficiently small—especially when it is comparable to or even smaller than in the Euclidean case—the present approximation is likely to provide an algebraic solution of the loop equations.
Furthermore, since the formal eigenvalue distribution $\rho_{M}(z)$ reproduces the perturbative expansion, if our approximation succeeds in reproducing this $\rho_{M}(z)$ with sufficiently small value of $g$, it strongly suggests that the approximation also captures a “physical” solution of the loop equations in such a small $g$ region. However, we emphasize once again that the perturbative expansion itself cannot be trusted for large values of $g$. Therefore, even if the optimization appears to work well in that regime, the resulting approximate values cannot immediately be regarded as physically meaningful. In the present paper we will not pursue the physical interpretation of the large-$g$ regime any further, and leave a detailed analysis to a forthcoming paper.
@@
Table 4: Bootstrap approximations for the Minkowski-type one-matrix model performed with $M=10$, $\Lambda=12$, $r_{n}=(0.8)^{n}$, and $g=\pm 1,-0.3,-0.1$. The quantities $a^{\text{formal}}$ and $\theta^{\text{formal}}$ are computed according to the formal solution given in Appendix B.
@@
With these considerations in mind, Figure 3 presents comparisons between $\rho_{M}^{(P)}(z)$ obtained from the bootstrap approximation and the formal solution $\rho_{M}(z)$ with g=1g=-1 (top panel) and g=0.3g=-0.3 (bottom panel). Both distributions are defined on line segments in the complex plane connecting $\pm e^{i\theta}a$. Here it should be emphasized that the parameters $\theta$ and $a$ are in general different for $\rho_{M}(z)$ and $\rho_{M}^{(P)}(z)$. To avoid confusion, we denote by $\theta$ and $a$ the parameters associated with $\rho_{M}(z)$, whose values are determined by (74), while $\theta^{(P)}$ and $a^{(P)}$ are those determined by the bootstrap approximation. With this notation, the supports can be written explicitly as
@@
Accordingly, $\rho_{M}(z)$ and $\rho_{M}^{(P)}(z)$ are defined on $\Gamma$ and $\Gamma^{(P)}$ as $\rho_{M}(e^{i\theta}\lambda)$ and $\rho_{M}^{(P)}(e^{i\theta^{(P)}}\lambda)$, respectively. When plotting the comparison between the two distributions, we evaluate $\rho_{M}(z)$ and $\rho_{M}^{(P)}(z)$ on their respective supports $\Gamma$ and $\Gamma^{(P)}$, and then display them on the same $\lambda$$\rho$ plane. As a consequence, the difference between the angles $\theta$ and $\theta^{(P)}$ cannot be read off directly from Figure 3 (however, as can be seen from Table 4, the numerical difference between these quantities is in fact extremely small). The numerical values of the relevant parameters are summarized in Table 4. Since the validity of the results cannot be guaranteed for large absolute values of $g$, we present here only four cases, $g=\pm 1,-0.3,-0.1$. However, we note that, based on our actual numerical computations, the optimization itself converges successfully even in the region $|g|>1$, with performance comparable to that for $|g|<1$.
@@
Finally, we comment on the phase transition of the Minkowski model. In the Euclidean model, there exists a critical value $g_{c}$ of $g$, beyond which a phase transition occurs and the theory breaks down. In the bootstrap approximation, this fact is presumably reflected in the failure of the optimization at $g=0.1$, even though the number of unknown variables is balanced with the number of equations. In contrast, in the Minkowski model, as far as we have investigated, no such behavior is observed, except for the instability around $g=0$. However, as we have repeatedly emphasized, the ansatz employed in this work essentially relies on the one-cut assumption, and in that sense it lacks generality. Even if one were able to remove this assumption and carry out the optimization, it is not clear whether the same behavior would still be observed. Therefore it would be premature at this stage to draw any definite conclusion regarding the absence of a phase transition.
In the case of the Euclidean one-matrix model, the existence of a critical coupling $g_{c}$ played an essential role in taking the double-scaling limit and establishing its equivalence with Liouville theory. In view of this physical importance, the possible existence of a phase transition in the Minkowski one-matrix model clearly deserves further investigation. For detailed calculations and discussions, see Appendix B.
@@ 5 Summary and Discussion
In numerical studies of matrix models, direct evaluations of matrix integrals using Monte Carlo methods have been extensively carried out, greatly advancing our understanding of theories primarily at finite $N$. On the other hand, within Monte Carlo approaches, taking the limit $N\to\infty$ is not possible in principle. From a theoretical point of view, however, the large-$N$ limit is expected to drastically simplify the structure of the system, while analytical calculations are in fact much more difficult at finite $N$, so this represents a rather curious inversion. Only recently has this situation begun to change, thanks to the development of matrix bootstrap methods based on semidefinite programming (SDP). Turning to our approach, we make more active use of the fact that large-$N$ matrix models are expected to possess a simple underlying structure, and translate this structure more directly into a numerical framework. One of the benefits of this approach is that the simplification due to the large-$N$ limit also extends to Minkowski-type matrix models, which makes it possible to apply the present method to the Minkowski-type one-matrix model under certain assumptions.
Concretely, in this work we have proposed a bootstrap approximation method based on the eigenvalue distribution $\rho(\lambda)$ and applied it to both Euclidean- and Minkowski-type one-matrix models. The method is designed to determine the (polynomially approximated) eigenvalue distribution $\rho^{(P)}(\lambda)$ and the values of the moments $w_{1},w_{2}$ self-consistently so as to satisfy the following three conditions: (i) there exists a probability distribution $\rho(\lambda)$ that generates the moments $w_{n}$; (ii) $\rho(\lambda)$ is a continuous function with finite support and can be well approximated by a finite-order polynomial; and (iii) the moments $w_{n}$ obey the loop equations. In practice, this is achieved by implementing a least-squares method that minimizes the objective function $F=\sum_{n}^{\Lambda}\lvert w_{n}-w_{n}^{(P)}\rvert^{2}$ . In addition, in order to apply this method to Minkowski-type one-matrix models, we have formally extended the notion of the eigenvalue distribution by relying on the large-$N$ factorization and the large-$N$ master-field conjecture. Under this assumption, the method can be applied to both Euclidean- and Minkowski-type models on the same footing. In fact, the numerical results agree extremely well with the exact solution in the Euclidean case and with the perturbative results in the Minkowski case.